2. Solve following questions (Inverse operator method) (10) y” + 2y+5y = xsin 2x
Answer #1

Jolution: Given equation is y +2y + 5y = x sin2x Let us considey y + 2y + 5y = xsin as (D+20+5) y = x 510 2x Auxilary equNow put 02-22 you non x sinex P -4+20+5 Se * sin 2x Rationalize the factor with 2D-1 2D-1 x sin 2x 40²,2 Yo = 20-1. x sin 2